### v1

# Enhancing Robot Safety via MLLM-Based Semantic Interpretation of Failure Data

Abstract
As robotic systems become increasingly integrated into real-world environments—ranging from autonomous vehicles to household assistants—they inevitably encounter diverse and unstructured scenarios that lead to failures. While such failures pose safety and reliability challenges, they also provide rich perceptual data for improving future performance. However, manually analyzing large-scale failure datasets is impractical. In this work, we present a method for automatically organizing large-scale robotic failure data into semantically meaningful clusters, enabling scalable learning from failure without human supervision. Our approach leverages the reasoning capabilities of Multimodal Large Language Models (MLLMs), trained on internet-scale data, to infer high-level failure causes from raw perceptual trajectories and discover interpretable structure within uncurated failure logs. These semantic clusters reveal latent patterns and hypothesized causes of failure, enabling scalable learning from experience. We demonstrate that the discovered failure modes can guide targeted data collection for policy refinement, accelerating iterative improvement in agent policies and overall safety. Additionally, we show that these semantic clusters can be employed for online failure detection, offering a lightweight yet powerful safeguard for real-time adaptation. We demonstrate that this framework enhances robot learning and robustness by transforming real-world failures into actionable and interpretable signals for adaptation.
1 Introduction
Autonomous systems—ranging from self-driving vehicles to household robots—are increasingly deployed in open, dynamic environments. In such unstructured settings, even state-of-the-art robotic systems are prone to failures due to unexpected interactions, unmodeled dynamics, and long-tail edge cases that deviate from training distributions. Traditional validation pipelines, often grounded in simulation or controlled testing, struggle to capture the full complexity of real-world deployment, leaving many failure modes undetected until operation.
A promising direction for improving robustness is to systematically learn from failures that occur during deployment. Robots naturally collect large volumes of perceptual data, including traces of both successful and failed interactions. These failure trajectories can provide valuable insights about the underlying conditions that led to safety violations, brittleness, or policy errors. However, manually curating and analyzing large-scale failure data is time-consuming and fundamentally unscalable.
In this work, we propose a framework for automatically organizing failure trajectories into semantically meaningful clusters, enabling scalable, unsupervised learning from failures. Our method leverages the reasoning capabilities of MLLMs—pretrained on internet-scale vision-language data—to infer high-level failure causes directly from raw observations. By applying prompted reasoning over sequences of perceptual inputs, we identify latent structure in uncurated failure logs, grouping them into interpretable categories described in natural language. Importantly, our framework operates in a completely unsupervised manner, obviating the need for costly human annotation while still isolating nontrivial keywords and cues associated with specific error causes.
The resulting semantic clusters offer multiple downstream benefits. First, they guide targeted data collection, enabling developers to focus training on critical or underrepresented failure modes. Second, these clusters can be integrated directly into online monitoring systems, providing a layer of semantic failure detection that acts as an early-warning mechanism for potential system safety violations during runtime. The semantic failure modes of the system can have many more downstream applications and use cases, e.g., targeted stress testing.
Our framework not only supports large-scale, structured analysis of failure data but also emphasizes interpretability – crucial for deployment in safety-critical domains. By revealing a structured understanding of failure causes in terms amenable to human interpretation, we offer actionable insights that can help stakeholders better assess the weaknesses of deployed systems and iterate more effectively.
Contributions.
To summarize, our key contributions are: (a) we propose a novel framework that uses MLLMs to cluster robotic failure data into semantically meaningful groups; (b) our method infers failure causes directly from raw perceptual sequences, eliminating manual annotation or supervision; (c) we demonstrate the effectiveness of our approach on large-scale failure datasets, as well as showcase its potential for supporting downstream applications, including targeted data collection and online failure detection; (d) our framework emphasizes interpretability, generating natural language summaries and keywords for each failure mode to support human-in-the-loop diagnosis.
2 Related Works
Semantic Clustering of Images with LLMs.
Recent research has demonstrated the effectiveness of Vision Language Models (VLMs) and Large Language Models (LLMs) for semantic grouping of images. For example, [24] clusters images using human-specified language criteria, while [26] removes the need for manual input by discovering clustering criteria directly from unstructured image collections. Other approaches like [16] highlight key semantic differences between image sets, and [29] and [37] focus on identifying important subpopulations or selecting semantically diverse subsets for training. Our work builds on this line of research by leveraging VLMs for semantic clustering; however, we focus specifically on failure trajectories in autonomous systems rather than static images, introducing a richer temporal and causal structure to the clustering task.
Text Clustering and Topic Modeling.
Recent topic modeling methods can be broadly grouped into three categories. First, extensions of classical topic modeling techniques [40, 44, 25, 8] like Latent Dirichlet Allocation (LDA) integrate word embeddings to enhance semantic representation [35, 38, 32, 27]. Second, fully embedding-based approaches [7, 11, 41, 6] that can leverage contextualized representations from pre-trained language models. Third, methods that separate the generation of the clusters from the generation of the topic representation, allowing for a flexible topic model [39, 2, 21]. BERTopic [21], extends this approach by incorporating a class-based variant of Term Frequency-Inverse Document Frequency (TF-IDF). Our work differs in that we use multimodal reasoning to discover semantic failure clusters directly from visual data and language-based explanations.
Failure Mining in Autonomous Systems.
Falsification has emerged as a prominent methodology for uncovering failures in autonomous systems. A variety of approaches [23, 18, 12, 19, 14, 45] have been proposed, wherein systems are tested in simulated environments with varied conditions designed to provoke failures. Such controlled testing allows researchers to identify the specific environmental parameters responsible for system breakdowns. However, while these methods effectively reveal failure scenarios tied to controlled variables, they fall short in capturing semantic failure modes of the system that require nuanced interpretation, requiring a human to manually go through those failure scenarios to identify failure modes of the system. In contrast, our approach leverages readily collected failure data to automatically discover semantic failure modes, bypassing manual human inspection.
Language-based Failure Reasoning in Robotics.
Recently, several studies have begun to integrate LLMs into robotic systems to generate language-based explanations of failures. For instance, [15] demonstrates the use of LLMs to generate explanations for failures in manipulation tasks, thereby improving system diagnostic capabilities. Similarly, [28] presents an LLM-based framework that not only produces informative failure explanations but also guides a language-based planner in correcting those failures for manipulation tasks. Our work extends these concepts to the domain of failure mode analysis, where the goal is to automatically cluster and interpret semantically rich descriptions from raw failure data, ultimately supporting downstream applications like enhanced monitoring and targeted data collection.
3 Problem Formulation: Finding Failure Clusters from Perception Recordings
We consider the problem of discovering semantic failure clusters from perception data collected during robotic failures. These clusters provide interpretable structure over uncurated failure logs and can be used for downstream tasks such as (1) targeted data collection to improve safety, and (2) online failure detection for real-time reliability.
Formally, we are given a dataset of sequences, each consisting of perceptual observations leading up to a failure:
where denotes the observation sequence for the -th failure case. The final observation corresponds to the failure event. For example, in autonomous driving, might be an image showing a rear-end collision.
Our goal is to construct a system that maps this dataset to a set of semantic clusters:
where each cluster is characterized by a natural language summary and a subset of sequences that share a common failure mode. These clusters reflect high-level failure themes derived from raw observations. For instance, a cluster might be: , in which case each sequence in corresponds to a failure where the autonomous car did not maintain a safe following distance from the vehicle in front. This formulation enables unsupervised discovery of interpretable failure modes directly from perceptual logs, providing a scalable mechanism for structuring real-world failure data and guiding robust, adaptive learning.
4 Method
4.1 Failure Cluster Discovery
Inferring the cause of failure in a robot trajectory is a complex task that requires understanding the robot’s environment, the agent’s actions, interactions with other agents, and their consequences. Doing so at scale across diverse failure episodes calls for automated systems capable of extracting and reasoning over semantic patterns in raw failure data. Our approach proceeds in three stages: (1) inferring failure reasons from perception sequences using an MLLM, (2) discovering semantic failure clusters via prompted reasoning over inferred causes, and (3) assigning each trajectory to one of the discovered clusters. The resulting clusters support both targeted retraining and online failure detection (Sec. 4.2).
Observation Downsampling.
To represent the failure event compactly, we downsample the tail of each sequence. Specifically, from each failure trajectory , we select the final frames, sampled from at a reduced frame rate. This balances the need to understand the temporal context of the failure with the limitations of the MLLM’s context window.
Step 1: Inferring Failure Reasons with MLLMs.
Each downsampled sequence is fed to an MLLM along with a structured prompt. The prompt first asks the model to describe the scene and agent behavior, then to infer a plausible cause of failure. We adopt a Chain-of-Thought (CoT) prompting strategy [43] to improve grounding and interpretability of the inferred reasons.
Step 2: Discovering Semantic Failure Clusters.
To uncover structure across failures, we provide all inferred failure reasons to a reasoning model that is tasked with clustering them into semantic failure modes. Each cluster is annotated with: a natural language name for the cluster, a short description of the failure type, keywords capturing representative situations in the cluster, and an estimated frequency of occurrence in the dataset. These annotations serve both interpretability and downstream usage in safety-critical settings.
Automated Prompt Ensemble and Cluster Aggregation.
Since LLM outputs are sensitive to prompt phrasing [46], we generate a prompt ensemble to improve robustness inspired by [34]. Given an initial prompt, an LLM proposes three other prompts using good practices for prompting. We use these ensemble of prompts to infer multiple failure clustering results. The resulting clusterings are merged by an aggregation model that consolidates overlapping clusters and unifies labels and descriptions into a single set of failure clusters . This approach generates a more comprehensive set of failure clusters by capturing diverse interpretations of the failure reasons and compiling them into one set of clusters.
Step 3: Assigning Trajectories to Failure Clusters.
Finally, each inferred failure reason is assigned to one of the discovered clusters by prompting a reasoning model with the cluster names , keywords, and descriptions generated in the previous step, and asking it to map each failure reason to one of the clusters. Sequences that do not match any cluster are flagged as outliers, which may indicate rare or complex failure modes.
4.2 Safety Enhancement
Discovered failure clusters support two key mechanisms for enhancing safety in autonomous systems.
Online Failure Detection.
Clusters, along with their associated keywords and explanations, form the foundation for semantic failure monitoring in our MLLM-based detection system. Rather than simple pattern matching, the system leverages this structured taxonomy of failure types and contexts to semantically interpret perceptual inputs. When the system recognizes potential failure conditions that align with known failure categories, it can raise warnings or trigger appropriate safeguard policies. This semantic understanding enables more nuanced and accurate failure detection as events unfold. The structured knowledge provided by clusters improves recognition of failures, enhancing the overall safety and reliability of the autonomous system.
Targeted Data Collection and Policy Refinement.
Semantic clusters enable targeted data augmentation by identifying classes of failures that require additional supervision. This allows practitioners to collect expert demonstrations or counterfactual data in these high-risk regimes and retrain the policy accordingly. Prior work has shown that specification-guided or failure-aware data collection improves robustness in previously unsafe scenarios [10, 22, 13, 36].
5 Experiments
To evaluate our framework for discovering semantic failure clusters, we present two case studies spanning distinct robotic domains. The first involves real-world dashcam footage of traffic incidents, and the second focuses on autonomous indoor navigation in office environments. In both cases, we aim to extract interpretable clusters of failure modes from raw perceptual sequences and demonstrate their utility for downstream tasks such as safety monitoring and targeted data collection.
We use Gemini 2.5 Pro [20] as the MLLM for failure reason inference, and OpenAI o4-mini [33] as the reasoning model for semantic cluster discovery, trajectory assignment, and failure monitoring. Prompt structures and other implementation details are provided in Appendix A.
5.1 Case Study 1: Real-World Car Crash Analysis from Dashcam Videos
We apply our framework to the Nexar driving dataset [30], which includes 1,500 dashcam videos (each approximately 40 seconds, 1280×720 resolution at 30 fps) of collisions or near-misses involving the ego vehicle. While these recordings are from human-driven vehicles, they serve as a proxy for autonomous driving failures in the absence of large-scale public AV failure datasets. However, our framework is directly applicable to autonomous vehicle logs as well.
5.1.1 Failure Cluster Discovery
Our system successfully discovers a diverse set of interpretable failure clusters from the driving dataset, as shown in Fig. 3. Qualitative inspection (Fig. 4) shows that these clusters correspond to meaningful and recurring traffic incident types, such as rear-end collisions, unsafe lane changes or intersection misjudgments.
Notably, the discovered clusters closely align with the U.S. DOT Volpe Center’s pre-crash typology [31], capturing most major failure types observed in real-world driving. This highlights our method’s ability to recover semantically grounded failure categories directly from unstructured video data in a way that aligns with expert-defined taxonomies.
Baselines and Results.
To evaluate the quality of the failure clusters discovered by our method, we compare against BERTopic [21], a state-of-the-art topic modeling method that combines transformer embeddings with unsupervised clustering and keyword extraction. We apply BERTopic to the same set of failure reasons produced in Step 1 of our pipeline to ensure a direct and fair comparison. We also evaluate a stronger variant, BERTopic-LLM, where a language model summarizes each discovered topic using representative keywords and samples.
Figures 5 and 6 illustrate the clusters and keywords produced by BERTopic and BERTopic-LLM, respectively. Standard BERTopic struggles to generate semantically coherent and interpretable clusters. Key failure modes — e.g., Rear-End Collisions— are diluted across vague or syntactically noisy topics (e.g., lead_vehicle_to_failed, white_the_lane_ego). BERTopic-LLM partially improves interpretability, recovering some valid modes such as Rear-end collision mechanics, but still includes vague or tautological categories such as Driving safety violations, which do not add diagnostic value or clearly differentiate failure modes. In contrast, our method consistently yields sharper, failure-relevant clusters (Fig. 3).
5.1.2 Failure Monitoring Leveraging Discovered Clusters
We evaluate runtime failure detection by prompting a multimodal reasoning model with recent image history and the learned failure clusters and reasoning about any possible near-future collision. Refer Appendix A.1.2 for more details.
Baselines.
We compare our method against VLM-Based Anomaly Detection (VLM-AD) methods [17]. These methods provide scene descriptions to an LLM and ask it to detect any possible anomalies in the current observation. We also compare against Leaderboard Method – the top-performing failure video classifier on the Kaggle Nexar challenge, based on fine-tuned VideoMAEv2-giant [42] Refer Appendix A.1.2 for details..
Results.
We compare all methods on a set of held-out driving trajectories. As shown in Table 1, our method consistently achieves the highest F1 score and outperforms baselines in both True Positive Rate (TPR) and False Negative Rate (FNR), showcasing its ability to detect the actual failures both robustly and reliably. These results also indicate that the system failures are not always the same as anomalies or out-of-distribution inputs; in-distribution scenarios, like those specified in our failure clusters, can also lead to system failures that are hard to capture with VLM-AD. We also test all methods on an unseen dashcam dataset of 200 trajectories (see Table 1 (Middle)). The proposed method demonstrates strong generalization. This suggests our method captures structured semantic patterns beyond dataset-specific cues, unlike the leaderboard classifier, which lacks generalizability.
Finally, we compare the lead failure detection time of different methods (Table 1 (Right)). Our method detects failures earlier than others, indicating a stronger ability to anticipate failures by correlating the scene observation with the failure clusters.
In-Distribution Trajectories | Out-of-Distribution Trajectories | ||||||||||
Method | TPR | TNR | FPR | FNR | F1 | TPR | TNR | FPR | FNR | F1 | Time |
Ours | 71.4 | 72.5 | 27.5 | 28.6 | 71.4 | 83.0 | 70.0 | 30.0 | 17.0 | 77.9 | 610 |
NoContext | 42.8 | 85.3 | 14.7 | 57.2 | 54.1 | 64.0 | 80.0 | 20.0 | 36.0 | 69.6 | 473.3 |
VLM-AD | 7.1 | 91.1 | 8.9 | 92.9 | 12.3 | 35.0 | 94.0 | 6.0 | 65.0 | 49.7 | 166.6 |
Leaderboard | 52.0 | 93.1 | 6.9 | 48.0 | 65.3 | 18.0 | 75.0 | 25.0 | 82.0 | 25.2 | 506.6 |
Ablations.
We also compare the performance of the proposed failure detection method on removing the failure cluster information from the prompt (NoContext). All other implementation details, such as processing a history of past observations and CoT reasoning, remain the same as our method. This results in significant performance degradation, highlighting the utility of failure clusters information in better detection across environments.
5.2 Case Study 2: Failure Discovery in Indoor Robot Navigation
We apply our framework to a vision-based ground robot, navigating unknown indoor office environments [4]. The robot uses a CNN-based policy with a model-based low-level controller. It receives RGB images, ego-velocities, and a goal position, and outputs acceleration commands for the robot. We record robot rollouts in the Stanford office environment [3] and extract front-view image sequences. Colliding trajectories comprise our failure dataset , which we use for clustering.
5.2.1 Failure Cluster Discovery
Our method discovers a set of interpretable failure clusters from the collision trajectories (Fig. 7). Notably, clusters such as walls and chairs were previously identified by manual inspection and reachability analysis in [10]. This validates our method’s ability to automatically recover known failure types and uncover new semantic patterns.
5.2.2 Targeted Data Collection and Policy Fine-Tuning for Enhanced Safety
Following Sec. 4.2, we use the discovered clusters to guide expert data collection in targeted regions of the environment. The robot policy is fine-tuned on this augmented dataset (details in Appendix A.2.3). The failure rate in sampled trajectories drops from 46% to 18%, demonstrating enhanced safety in previously failure-prone situations. This forms a closed-loop pipeline of failure discovery, targeted intervention, and policy refinement for continuously enhancing system safety.
5.2.3 Failure Monitoring
We use the same runtime monitoring approach as in Case Study 1. The monitor reasons over the scene, past trajectory, and known failure clusters to preemptively identify any potential failure. As evident in Fig. 8, due to the confined and cluttered nature of the indoor environment, the failure monitor can easily misinterpret static background elements with prior failure contexts, leading to a high false positive rate. To mitigate this, we introduce a simple temporal consistency rule: a failure is flagged only if it persists for three consecutive frames. This helps reduce conservativeness while preserving responsiveness. Refer Appendix A.2.2 for details.
Baselines.
We compare against VLM-AD, where we adapt the prompt from [17] by giving examples relevant to an indoor office environment in the prompt. We additionally compare against ENet-BC, a vision-based runtime failure monitor based on EfficientNet-B0, trained on labeled collision data from the same environment [10]. Note that, unlike our method and VLM-AD, ENet-BC requires environment-specific training and does not generalize to unseen layouts (details in Appendix A.2.2).
Results.
On an in-distribution (IID) test set of 326 trajectories, our method outperforms all LLM-based baselines in F1 score (Table 2, Left). ENet-BC achieves similar performance to the proposed method, as expected given its environment-specific training.
In-Distribution Trajectories | Out-of-Distribution Trajectories | ||||||||||
Method | TPR | TNR | FPR | FNR | F1 | TPR | TNR | FPR | FNR | F1 | Time |
Ours | 65.0 | 99.0 | 1.0 | 35.0 | 77.2 | 67.6 | 86.2 | 13.8 | 32.4 | 50.0 | 1.21 |
NoContext | 51.7 | 99.6 | 0.4 | 48.3 | 67.37 | 45.9 | 89.0 | 10.1 | 54.1 | 40.5 | 0.76 |
VLM-AD | 83.3 | 50.0 | 50.0 | 16.7 | 40.0 | 62.2 | 60.1 | 39.9 | 37.8 | 27.2 | 1.38 |
ENet-BC | 65.0 | 100.0 | 0.0 | 35.0 | 78.8 | 100.0 | 6.3 | 93.7 | 0.0 | 22.4 | 1.01 |
To test generalization, we evaluate on an out-of-distribution (OOD) set of 300 trajectories from a different building. The performance of all methods degrades as expected, but the proposed method maintains the highest F1 score, while ENet-BC fails entirely to generalize. This again demonstrates the generalization capabilities of our method.
Our method also detects failures earlier on average, highlighting its ability to reason about impending collisions before impact. Full metrics are in Table 2. Although VLM-AD has a slightly higher average time, it is because it fails entirely and detects everything as anomalous, evident from its high FPR.
Engaging the Safeguard Policy
We integrate our failure monitor with a reactive safeguard controller that activates upon a failure detection. Fig. 8 shows an example where the nominal policy leads to a collision, while the monitor detects the failure and invokes the safeguard controller, enabling successful recovery.
6 Conclusion and Limitations
We present a closed-loop framework for interpretable failure analysis in autonomous systems. Our method automatically discovers semantically meaningful failure modes from perception recordings without supervision, distilling them into human-understandable clusters annotated with natural language. These clusters support targeted data collection, policy fine-tuning, and semantic failure detection—enabling scalable and continuous safety improvement. By leveraging the reasoning capabilities of MLLMs, our approach provides a foundation for understanding, organizing, and responding to failures in complex, unstructured environments.
While our framework enables interpretable and actionable failure analysis, it has several limitations. First, there is no canonical “correct” way to cluster failure trajectories. Different clustering strategies may highlight different operational insights. Future work could integrate formal techniques like STPA (Systems-Theoretic Process Analysis) or FRAM (Functional Resonance Analysis Method) to complement unsupervised clustering. Second, MLLMs can generate plausible-sounding but incorrect failure explanations, or miss the true cause entirely. Future work could incorporate simulation-based counterfactual testing or causal world models to verify and refine inferred explanations. Finally, while we demonstrate results on datasets of 700 trajectories, our pipeline can scale to larger datasets and temporally extended events with appropriate batching and model throughput optimization.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge research support from DARPA Assured Neuro Symbolic Learning and Reasoning (ANSR) program and by the NSF CAREER program (2240163).
Appendix A Technical Appendices and Supplementary Material
The appendix offers additional details, such as prompts, about our implementation. Sections A.1 and A.2 provide the details for dashcam driving datasets and indoor ground navigation, respectively.
A.1 Case Study 1: Real-World Car Crash Analysis from Dashcam Videos
Section A.1.1 contains all the details of discovering the failure clusters, and Section A.1.2 lists all the details about runtime monitoring for the driving case study. The grey boxes contain prompts.
A.1.1 Failure Cluster Discovery
Step 1: Inferring Failure Reasons with MLLMs.
Step 2: Discovering Semantic Failure Clusters.
We generate an ensemble of prompts from an LLM (OpenAI o4-mini) to infer multiple failure clustering results, which are then aggregated into the final list of clusters. Here are the three prompts generated by the LLM:
Prompt 1:
Prompt 2:
Prompt 3:
Each of these prompts is appended with a list of all the failure reasons from the previous step. Finally, we query the LLM again to consolidate these individual lists and generate the final clusters.
Step 3: Assigning Trajectories to Failure Clusters.
We compare the distribution of discovered clusters to their actual assignments across trajectories (Fig. 9). Most frequencies align closely. The largest discrepancies appear for the visibility and sudden intrusion clusters. Visibility failures show higher reported than assigned frequency in single-label settings, but match under multi-label assignment—suggesting co-occurrence with other failures. Similarly, sudden intrusions is under-assigned due to ambiguity with intersection or cut-in cases.
A.1.2 Failure Monitoring Leveraging Discovered Clusters
As described in Sec. 4.2, we directly use the obtained failure clusters to safeguard the system through runtime monitoring. The reasoning model takes a history of past image observations as input and tries to reason about any possible near-future collision while leveraging the knowledge about the most common failures of that system from given clusters in a Chain-of-Thought (CoT) manner. Adding the cluster information in the runtime monitor’s prompt helps in zero-shot failure detection for timely executing a safe action, and generalizing across different agents and environments without requiring any specialized training. Here is the prompt:
The prompt first outlines the main task for the LLM to act as a runtime failure monitor and detect if the system is currently SAFE or UNSAFE based on the observation history. Further, it provides more information about the system, such as what is meant by a failure here and the frequency of observations. Then, we ask it to sequentially reason about the possible future motion of the system, the surrounding environment, and any possible risk of failure given the most common failure modes of that system through clusters. list of clusters contains all the generated clusters along with their keywords and descriptions.
We report the failure detection metrics on both In-Distribution (IID) Trajectories and Out-of-Distribution Trajectories (OOD) in Table 1. For IID testing, we took a held-out set of 200 trajectories from the same Nexar dataset, which we used for clustering analysis. For OOD testing, we took another open-source dashcam dataset [5] containing ego car crash videos.
Fig. 10 shows a few examples for True Positives and False Negatives predictions from our method. Rows 1,2 are successfully detected as unsafe. Rows 3,4 are failures, but are detected as safe because the failure event is not evident in those images. Note that the last image in the sequence is taken 1 sec before the collision, and since car driving is a high-speed system, the actual failure sometimes starts even after that, which leads to false safe predictions.
Baselines.
VLM-AD is an LLM-based few-shot runtime anomaly detection method, where the current image observation is first processed through a scene descriptor, such as an object detector, which further feeds into an LLM. The LLM is asked to find any possible anomalous scenario in the current scene, by providing it with a description of some Normal and Anomalous examples in the prompt. For the driving case study, we directly take the prompt from [17] and use the same reasoning LLM as our method to prevent any model-specific performance difference and perform a fair comparison of the methods. It should also be noted that these methods require manually writing few-shot examples in the prompt to provide contextual information about the system, which in our method is automated with the generated clusters.
For the Leaderboard-Method baseline, since the Nexar dataset is also hosted as a Crash Prediction Challenge on Kaggle, we take the best publicly available method [1] as per leaderboard scores to compare with our method. It takes a pre-trained VideoMAEv2-giant model as the backbone and trains a Linear binary classification MLP layer to predict crash. VideoMAEv2-giant [42] is trained on a large amount of video datasets, such as action recognition datasets, sports videos, instructional videos, etc. We tested their trained model on our test dataset for both the IID and OOD cases. Being a classical vision-based classification approach, this method requires environment-specific training and fails to generalize.
A.2 Case Study 2: Failure Discovery in Indoor Robot Navigation
Section A.2.1 contains all the details of discovering the failure clusters, and Section A.2.2 and Section A.2.3 contain all the details about runtime monitoring and retraining for the indoor robot navigation, respectively. The grey boxes contain prompts.
A.2.1 Failure Cluster Discovery
Step 1: Inferring Failure Reasons with MLLMs.
We randomly sample initial states in the environment and trajectory rollouts with a fixed goal location. For each rollout, we record the front-view image observations and a binary label denoting Failure or Success. Further, we filter out the failure trajectories for our purpose. The following clustering analysis is performed on 228 failure trajectories.
Step 2: Discovering Semantic Failure Clusters.
Here is the ensemble of prompts given by the LLM. Each of these prompts is appended with a list of all the failure reasons from the previous step.
Prompt 1:
Prompt 2:
Prompt 2:
Prompt 3:
Step 3: Assigning Trajectories to Failure Clusters.
We compare the distribution of discovered clusters to their actual assignments across trajectories (Fig. 9). Most frequencies align closely. The largest discrepancies appear for the Narrow-Gap/Clearance Misjudgment and Low-Height Clutter & Small Floor Obstacles clusters.
They show higher reported than assigned frequency in single-label settings, but match under multi-label assignment–suggesting co-occurrence with other failures. Similarly, Bins & Waste Receptacles is under-assigned due to ambiguity with Low–Height Clutter & Small Floor Obstacles or Box--Like Equipment & Carts.
A.2.2 Failure Monitoring Leveraging Discovered Clusters
We perform runtime failure detection in the same way as described in A.1.2 for the indoor navigation robot. The prompt is tailored to an indoor robot with its list of clusters. Here is the detailed prompt:
We report the failure detection metrics on both IID and OOD test sets 2. The IID test set was taken from the same environment, but the OOD test set was taken from a different building with very different semantics, but an indoor office environment in the dataset.
Fig. 12 shows a few examples for True Positives and False Negatives predictions from our method. Rows 1,2 are successfully detected as unsafe. Rows 3,4 are failures, but are detected as safe because the monitor is getting confused seeing some free space ahead, and the temporal consistency rule gives false safe outputs.
Baselines.
For the VLM-AD baseline, we adapt the prompt from [17] for an indoor robot and test it in the same way as described in A.1.2. Here is the prompt:
For the ENet-BC baseline, we follow [9] and first compute the Backward Reach-Avoid Tube (BRAT) for our environment. Further, we collect a dataset of 3000 images with 1500 safe and 1500 unsafe images labeled using the BRAT and follow [22] to train the EfficientNet-B0 model for binary classification. Similar to the Leaderboard method in the driving example, it is based on environment-specific training, and there’s no opportunity for generalization across environments, where LLM-based methods shine.
A.2.3 Targeted Data Collection and Policy Fine-Tuning for Enhanced Safety
For targeted data collection around the mined failure modes of the robot, we identify regions in the environment corresponding to each cluster. In each failure region, for some starting states and a goal state, we use a Model Predictive Control (MPC) scheme to find a sequence of dynamically feasible waypoints [4]. The collected dataset around all the failure modes comprised approximately 40K training samples, which were then augmented with the original training data. Lastly, we fine-tuned the trained checkpoint for only 20 iterations on the augmented dataset and compared the performance on the previous policy and the retrained policy by rolling out trajectories for 50 initial states in the same environment.
Fig. 13 shows a few examples from the targeted dataset.
Generated clusters with keywords and their descriptions for driving dashcam datasets.
Generated clusters with keywords and their descriptions for indoor robot navigation.